OpenStudy (kkutie7):

If 0.4506 g of KHP standard were titrated with NaOH and the endpoint was reached at 27.96 mL, what is the concentration (molarity) of the NaOH solution?

OpenStudy (nincompoop):

let me see you solve it first

OpenStudy (kkutie7):

I can't that's the problem... I'm not sure where to start.

OpenStudy (nincompoop):

what is titration? what is endpoint?

OpenStudy (nincompoop):

you must know something to begin with

OpenStudy (cuanchi):

KHP, stand for potassium hydrogen phthalate with a MM 204.22 g/mol and react one mole to one mole with NaOH (~40g/mol)

OpenStudy (nincompoop):

what is equilibrium? something to begin with

OpenStudy (kkutie7):

\[.4506g KHP*\frac{ 1mole KHP }{ 71.0799g KHP }*\frac{ 1moleNaOH }{ 1 mole KHP }*\] \[\frac{1}{.02796L}\] does this look even a bit right?

OpenStudy (kkutie7):

I figure endpoint should mean 1:1 moler ratio

OpenStudy (cuanchi):

The MM of the KHP is NOT 71.0799 g/mol, it is 204.22 g/mol (C8H5KO4)

OpenStudy (kkutie7):

my mistake.... but how about the rest of my problem?

OpenStudy (cuanchi):

http://en.wikipedia.org/wiki/Potassium_hydrogen_phthalate the rest is fine

OpenStudy (nincompoop):

first write the balanced chemical equation

OpenStudy (nincompoop):

it is asking for molarity so your dimensional analysis must be able to indicate molarity (solute per liter) \[M = \frac{moles_{solute}}{Liters_{solution}}\]

OpenStudy (kkutie7):

was what I did wrong?

OpenStudy (kkutie7):

\[C_8H_5KO_4+NaOH \rightarrow KNaC_8H_4O_4 +h_2O\] is this good enough?

OpenStudy (kkutie7):

\[.4506g KHP*\frac{ 1mole KHP }{ 71.0799g KHP }*\frac{ 1moleNaOH }{ 1 mole KHP }*\] \[\frac{1}{.02796L}\] whats wrong with is?

OpenStudy (kkutie7):

*this

OpenStudy (nincompoop):

submit it and see

OpenStudy (kkutie7):

=/ I'm not working with sapling or something like that. I can't just submit it for a check.

OpenStudy (nincompoop):

that is how it is during exam you don't get to ask your professor if it is correct before you submit

OpenStudy (nincompoop):

why aren't you confident with your solution is the question

OpenStudy (cuanchi):

Change the molecular mass from 71 to 204 !!!

OpenStudy (nincompoop):

LAUGHING OUT LOUD I thought you corrected that one already @Cuanchi

OpenStudy (nincompoop):

this is why I asked to write the balanced chemical equation

OpenStudy (nincompoop):

did you use the masses of potassium, hydrogen and phosphorus?

OpenStudy (cuanchi):

Many students commit this mistake

OpenStudy (nincompoop):

your analysis is correct by the way I was adamant about the balanced equation so you can spot where the mistake was